Two firms compete in a market to sell a homogeneous product with inverse demand function P = 600 – 6Q. Each firm produces at a constant marginal cost of $300 and has no fixed costs. Use this information to compare the output levels and profits in settings characterized by Cournot, Stackelberg, Bertrand, and collusive behavior. Please show steps.
Answer #1

Cournot 0 P=600-60 mc, = MC2 = 300 The P, CO-GO, = 6000,-60²-600-3000 600-1201-60,-3ou no 01 = 300,250250202 similaily for- s